OpenStudy (anonymous):

There are 70 teachers in a school. Of the 38 male teachers, 8 teach Maths It is known that 2/7 of all teachers teach Maths. Find the probability that a teacher selected at random is. (a) a male teacher or a teacher who teaches Maths. (b) a female teacher or a teacher who teaches Maths. (c) a female teacher or a teacher who does not teach Maths.

OpenStudy (tkhunny):

Draw a 2x2 grid with total. Fill in the blanks.

OpenStudy (anonymous):



OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

thanks for your help. However, the answer is 5/7 ..

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@ajprincess

OpenStudy (anonymous):

Your 2 x 2 grid looks good. Which question is still unanswered?

OpenStudy (anonymous):

part a . haha. i cannot calculate the correct answer.

OpenStudy (anonymous):

I think it is all but the square that is female non-maths, as the other squares are included in the requirements. Thus, 58/70?

OpenStudy (anonymous):

the answer is 5/7

OpenStudy (anonymous):

Which must come from 50/70. Or [8+12+30 ]/70

OpenStudy (anonymous):

where comes 30?

OpenStudy (anonymous):

Try this males = 8, non-maths males =30, but we have already counted 8 males in that, so those two are 8 + (30-8) = 30 and then 20 female non-maths gives 50 and the total denominator is 70.

OpenStudy (anonymous):

but the question is asking "a male teacher or a teacher who teaches Maths"....

OpenStudy (anonymous):

Yes, but we cannot double-count the males who do teach maths. I think I am right, but I am not sure.

OpenStudy (anonymous):

However, i think 50 is non-maths teachers but not maths teachers....

OpenStudy (anonymous):

i can ask my maths teacher tomorrow. let move onto part b first

OpenStudy (anonymous):

Try: 1 minus all excluded. I think that is the easier way to see it and get 50/70.

OpenStudy (anonymous):

this method is called complementary events, isn't it? So difficult..

OpenStudy (anonymous):

Not sure of terminology. Yes, difficult.

OpenStudy (anonymous):

and part b is similar to part a...

OpenStudy (anonymous):

Yes. "OR" is the key. Means we cannot just multiply two independent probabilities as we can with "AND." OR means "either or both."

OpenStudy (anonymous):

and part c ? 32/70+30/70=31/35?

OpenStudy (anonymous):

i think the formula doesn't make sense..

OpenStudy (anonymous):

ahh, i know how to do now! for part a, "a male teacher or a teacher who teaches Maths" doesn't mean the male teacher teaches Maths! Therefore, a male teacher=38 a teacher who teaches Maths=12 (which is the number of female teachers as you said it cannot be double-count) So, (38+12)/70=5/7

OpenStudy (anonymous):

and for part b, a female teacher=32; a teacher who teaches Maths=8(which is the number of male teachers who teach Maths) Therefore, 40/70=4/7 .......does it make sense?

OpenStudy (anonymous):

"a male teacher or a teacher who teaches math" includes all males, and all those who teach math, except that we have already counted 8 of them as males.

OpenStudy (anonymous):

all males means male teachers who teach non-Maths and Maths??

OpenStudy (anonymous):

yes, all males means all of them regardless of what they teach.

OpenStudy (anonymous):

oh i see. No wonder why i confused these questions. And three parts i got the correct answer. Thanks for your help!!:) do you know where does the permutation and combination concepts post in openstudy?

OpenStudy (anonymous):

No, but I know that permutations consider order crucial, and the number of permutations of n things is n!. Combinations ignore order, and the number of combinations of n things taken r at a time is n!/(r!)(n-r)! These are the central concepts and equations.

OpenStudy (anonymous):

does "ignore order" mean without repetition? thank you so much! I must go to bed now as my time here is 11 pm already. Goodnight and million thanks:)

OpenStudy (anonymous):

ignore order means the items are considered identical

OpenStudy (anonymous):

I am in new York, here are you?

OpenStudy (anonymous):

good night